OpenStudy (anonymous):

How can one type symbols on here such as exponents or radical signs , if theeyre not on the keypad?

OpenStudy (amistre64):

\*[\sqrt{4x^7} \*] <-- remove the *s to format \[\sqrt{4x^7} \]

OpenStudy (amistre64):

\*[\sqrt[3]{27}\*] \[\sqrt[3]{27}\]

OpenStudy (amistre64):

the other method is to hunt thru the equation button